OpenStudy (anonymous):

x/2+2/3=5/6 how to get the value of X

OpenStudy (ybarrap):

x/2+2/3 = 5/6 Subtract 2/3 from both sides: x/2 = 1/6 Divide both sides by 1/2: x = 1/3

OpenStudy (anonymous):

thanks